OpenStudy (anonymous):

given sec(theta)= 5... find cot(90-theta)

OpenStudy (asnaseer):

do you know what sec(theta) represents in terms of cos(theta)?

OpenStudy (anonymous):

yesh.. it represents its recioprical

OpenStudy (anonymous):

pardon the spelling

OpenStudy (asnaseer):

:) thats ok - we all make mishtakes :D

OpenStudy (anonymous):

asnaseer why did u report me

OpenStudy (anonymous):

wlbrt316 were busy in here..

OpenStudy (asnaseer):

ok, so you know:\[\sec(\theta)=\frac{1}{\cos(\theta)}\]which means you now know:\[\cos(\theta)=\frac{1}{\sec(\theta)}=\frac{1}{5}\]

OpenStudy (anonymous):

great so now we know cos(theta). but i dont see how that helps us find cot(90-theta)

OpenStudy (asnaseer):